# Low Mass Naked Singularities from Dark Core Collapse

### v1

@@ 4.1 DM particles cloud modeled as a perfect fluid or inhomogeneous dust
$T^{\alpha}_{\beta}={\rm diag}\left(-\rho,p,p,p\right),$ (16)
@@
$\rho=\frac{F^{{}^{\prime}}}{R^{{}^{\prime}}R^{2}}$ (18)
@@
$R(t,r)=\left(r^{\frac{3}{2}}-\frac{3}{2}\sqrt{F(r)}t\right)^{\frac{2}{3}},$ (20)
@@
$t(r,s)=\frac{2r^{3/2}}{3\sqrt{F(r)}}\left(1-s^{3/2}\right)$ (21)
@@
$t_{v}(r)=\frac{2r^{\frac{3}{2}}}{3\sqrt{F(r)}}$ (22)
by substituting $R(t_{v},r)=0$ in Eq. (20) or $s=0$ in Eq. (21). This, in fact, gives the singularity curve which depicts the spacetime singularity formation as the collapse endstate. Note that the collapsing DM fluid spacetime (Eq. 17) is matched smoothly with the exterior Schwarzschild spacetime. However, the singularity formed in the DM dust collapse scenario is event-like singularity. This means that the singularity initially becomes visible in principle, as the event horizon does not form. In a later stage, the singularity gets covered within the event horizon, provided that the energy conditions ensuring the positivity of mass and energy density are satisfied. Thus, the conclusion is, the formation of a tiny endoparasitic BH could not be avoided eventually for an inhomogeneous dust collapse. The similar situation could arise, even if the DM cloud is modeled as a perfect fluid, and, hence, the conclusion remains same.
@@ 4.2 DM particles cloud modeled as an anisotropic fluid
$ds^{2}=-e^{2\nu}dt^{2}+\frac{R^{\prime 2}}{H}dr^{2}+R^{2}d\Omega^{2}\;$ (23)
@@
Eqs. (23) and (24) reveal that one obtains six unknowns ($\rho,p_{\theta},\nu,H,R$ and the mass function $F$ for the DM cloud) that appear due to the integration of the Einstein equations. As there are four Einstein equations, we are free to choose two functions. The free functions are chosen to be the mass function $F(r)$, which is the initial mass that is conserved for the DM cloud, and the tangential pressure of it. Although $F(r)$ is shown to be independent of time [10], the tangential pressure depends on $r$ and $t$ via [10] $v(r,t)$ as $p_{\theta}\equiv p_{\theta}(r,v)$. A global choice for the pressure at all times fully fixes the evolution of the system.
@@
$F(r)=M_{\rm e}r^{3},\quad v_{\rm e}(r)=r^{\alpha},\quad F(R)=M_{\rm e}R^{\frac{3}{\alpha+1}},$ (25)
@@
$ds^{2}_{e}=-(1-M_{\rm e})\left(\frac{R}{R_{\rm b}}\right)^{\frac{M_{\rm e}}{1-M_{\rm e}}}dt^{2}+\frac{dR^{2}}{1-M_{\rm e}}+R^{2}d\Omega^{2}\;,$ (26)
@@
$\frac{F}{R_{\rm e}}=1-H_{\rm e}=\frac{4p_{\theta\rm e}}{\rho_{\rm e}+4p_{\theta\rm e}}\;.$ (27)
@@ 5 Transmutation timescale of the host neutron star
In Sec. 4, we show that the endoparasitic collapsed object could be a BH (see Sec. 4.1) or a naked singularity (see Sec. 4.2) depending on the initial conditions. However, after its formation it will start to accrete matter immediately from its host neutron star with the spherical Bondi accretion rate:
$\frac{dm}{dt}\Bigg{|}_{\rm acc}=k_{1}m^{2}\,,$ (28)
@@
Figure 1: Mass ($M_{0}$) of the endoparasitic collapsed object (BH or naked singularity) formed inside a neutron star of $M_{\rm h}=1.4M_{\odot}$ at $t=t_{0}$ versus the mass ($m_{\chi}$) of DM particles. This figure explains that $M_{\rm h}>>M_{0}$. See Sec. 5 for details.
Now, by integrating Eq. (28):
@@
$\Rightarrow t_{\rm d}-t_{0}$ $=$ $\frac{1}{k_{1}M_{0}}\left[1-\frac{M_{0}}{M_{\rm h}}\right],$ (29)
@@
$\Rightarrow t_{\rm d}$$\approx$ $t_{0}+\frac{c_{\rm s}^{3}R_{\rm h}^{3}}{3G^{2}M_{\rm h}M_{0}}\equiv t_{0}+t_{\rm acc}$ (30)
where, $t_{\rm acc}$ is the accretion timescale after the formation of endoparasitic collapsed object. To obtain Eq. (30) from Eq. (29), we neglect the term $M_{0}/M_{\rm h}$ as $M_{\rm h}>>M_{0}$ which is clear from Fig. 1. The solid magenta and dashed purple curves of Fig. 1 indicate the mass ($M_{0}$) of the newly-formed endoparasitic collapsed object inside a typical neutron star of $M_{\rm h}=1.4M_{\odot}$ by collapsing of particles, respectively. Note that the Bondi (spherical) accretion rate (Eq. 28) should be same [49] for a BH and a naked singularity. Since this should also be true for Eq. (30), we only need to know $M_{\rm h}$ and the mass of the endoparasitic collapsed object $M_{0}$ (see Eq. 13) to calculate $t_{\rm d}$, regardless of its nature.
@@
Figure 2: The variation of transmutation timescale ($t_{\rm d}$) of a neutron star of $M_{\rm h}=1.4M_{\odot}$ and $R_{\rm h}=10$ km as a function of the mass ($m_{\chi}$) of non-self-annihilating DM particles. Panels (a)-(b) correspond to two different DM density profiles with a fixed DM particle-nucleon scattering cross section, i.e., $\sigma>10^{-44}$ cm2. The black dotted horizontal line represents the Galactic age ($t_{\rm Ga}$). This figure shows that there is a non-zero possibility of transmutation of the neutron star into a collapsed object (BH or naked singularity depending on the initial conditions) in a reasonable time for a wide range of $m_{\chi}$ values, depending on the value of DM density profile ($\rho_{\rm DM}$). See Sec. 5 for details.
@@
Figure 3: Constraints on the DM particle-neutron scattering cross section $\sigma$, as a function of mass ($m_{\chi}$) of the DM particles for the possible transmutation of a typical neutron star, if it is located at a distance $D$ with the specified $\rho_{\rm DM}$ value mentioned in the caption of each panel. The parameter space located on the right side of the red/blue-dashed curve indicates a transmutation for $t_{\rm d}<t_{\rm Ga}$. On the other hand, the left side of the red/blue-dashed curve suggests that the transmutation may occur in the future. See Sec. 5 for details.
@@
For the purpose of demonstration to calculate $t_{\rm d}$, let us assume that a DM particle is captured after a single scattering, and, hence, we consider here the non-self-annihilating DM particles with a mass range: $1$ GeV $<m_{\chi}<10^{6}$ GeV [26]. If a typical neutron star of $M_{\rm h}\sim 1.4M_{\odot}$ and $R_{\rm h}\sim 10$ km captures the DM particles in the above mass range and forms an endoparasitic BH/naked singularity in its core, the entire neutron star could transmute into a BH/naked singularity depending on the initial condition of collapse as discussed in Sec. 4. Now, if the transmutation time ($t_{\rm d}$) of a neutron star satisfies the following condition: $t_{\rm d}<t_{\rm Ga}$ (where $t_{\rm Ga}\sim 1.36\times 10^{10}$ yr is the Galactic age), the neutron star could have already transmuted. Note that $t_{\rm d}$ depends on the DM density ($\rho_{\rm DM}$) profile (see Eq. 7, Eq. 14 and Eq. 30) as well. Thus, for the purpose of demonstration, here we consider the most commonly used Navarro-Frenk-White (NFW) DM density profile [50],
$\rho_{\rm DM}(D)$ $=$ $\rho_{0\rm DM}~{}(D/D_{s})^{-\alpha}(1+D/D_{s})^{-3+\alpha}$ (31)
@@
where, $D$ is the distance from the Galactic center. We consider the local DM density $\rho_{0\rm DM}=0.4$ GeV.cm-3 following [50, 45], the scale radius $D_{s}=20$ kpc following [31] with $\alpha=1$ [31]. Thus, $\rho_{\rm DM}\sim 7\times 10^{4}$ GeV.cm-3 [31] is obtained at a distance $D\sim 0.1$ pc inside the Galactic center, and $\rho_{\rm DM}\sim 0.4$ GeV.cm-3 is for the Galactic disk ($D\sim 10$ kpc) [51]. Substituting all the values in Eq. (30), one can plot $t_{\rm d}$ versus $m_{\chi}$ for the DM particle-neutron cross section $\sigma>10^{-44}$ cm-2 for the two different values of DM density.
@@
Panel (a) of Fig. 2 shows that the transmutation of a neutron star could not be possible in the Galactic disk for the fermionic DM particles, whereas it could be possible for the bosonic DM particles. Panel (b) corresponds to the Galactic bulge, and it shows that the transmutation could be possible for both the fermionic and bosonic DM particles. However, comparing both the panels we can see that the possibility of transmutation within $t_{\rm d}\leq t_{\rm Ga}$ decreases with decrement of $\rho_{\rm DM}$. Note that Fig. 2 stands for $\sigma>10^{-44}$ cm-2 only, and, thus it can be considered just as an example of how $t_{\rm d}$ varies with $m_{\chi}$ for three different $\rho_{\rm DM}$ values with a fixed value of $\sigma$. One cannot constrain the $\sigma-m_{\chi}$ parameter space from Fig. 2. Panels (a)-(b) of Fig. 3, in fact, help to constrain the non-self-annihilating DM-neutron scattering cross section $\sigma$, as a function of $m_{\chi}$ for the possible transmutation within $t_{\rm d}\leq t_{\rm Ga}$, if the neutron star is located in the Galactic disk and Galactic bulge, respectively. The parameter space located on the right side of the red (blue-dashed) curve indicates a transmutation for $t_{\rm d}\leq t_{\rm Ga}$. On the other hand, the left side of the red (blue-dashed) curve suggests that the transmutation may occur in the future.
@@ 6 End state of the transmuted object
Now, the question is, what would be the end state of the transmuted object. The tiny endoparasitic BH formed (as discussed in Sec. 4.1) due to the collapse of a perfect DM fluid or dust should transmute the host into a BH. The transmutation process should be as discussed in various papers (e.g., [25, 27]). If the host neutron star has a non-zero spin (e.g., a pulsar), the transmutation process could be evolved as described in [52]. However, the end state of the transmuted object would be a BH, and not a naked singularity as shown in [53].
When an endoparasitic naked singularity (as discussed in Sec. 4.2) is formed due to an anisotropic DM fluid collapse, it starts to accrete matter from its host and eventually the entire host would transmute into a naked singularity. As an exactly non-spinning object may not exist in the Universe [54], the host neutron star should have a non-zero spin parameter $a_{\rm h}$ $\equiv J_{\rm h}/M_{\rm h}$, where $J_{\rm h}$ is the angular momentum of the host. Since the spin frequency ($\omega$) of both the neutron star and the endoparasitic naked singularity during its formation (i.e., at ${t\rightarrow t_{0}}$) should be the same (as $M_{0}<<M_{\rm h}$) [53], the endoparasitic naked singularity should also have a non-zero spin parameter. Let us denote that spin parameter of the endoparasitic naked singularity as $a_{0}$ at time $t\rightarrow t_{0}$. However, if one calculates the numerical value of $a_{0}/M_{0}$, it comes as negligible small (as $M_{0}<<M_{\rm h}$), and, hence, one can consider it as $a_{0}\rightarrow 0$ for all the practical purposes. During accretion from its host, the value of spin paramater starts to increase from $a_{0}$ and could eventually acquire $\sim a_{\rm h}$ [53] at the end of the transmutation process (i.e., $t=t_{\rm d}$), if the dynamical ejecta is small [52, 53]. If the dynamical ejecta is large [53] during accretion by the endoparasitic naked singularity, the final transmutation could give birth to a sub-solar-mass or near-solar mass slowly-spinning JMN-1 naked singularity. Computation of the evolution of such an endoparasitic naked singularity, which is a complex and entirely unexplored field, will be considered in future research works.
@@
$-2a\sin^{2}\theta\left[\left(\frac{R}{R_{\rm b}}\right)^{\frac{M_{\rm e}}{2(1-M_{\rm e})}}-(1-M_{\rm e})\left(\frac{R}{R_{\rm b}}\right)^{\frac{M_{\rm e}}{(1-M_{\rm e})}}\right]dtd\phi$ (32)
where, $a$ is the spin parameter of this spacetime. The above slowly-spinning JMN-1 spacetime (Eq. 32) can be matched with the exterior slowly-spinning Kerr spacetime at $R=R_{\rm b}$ [55],
@@
where $2M=M_{\rm e}R_{\rm b}$ is the Schwarzschild mass of the slowly-spinning ($a/M<<1$) naked singularity. Eq. (32) reduces to Eq. (26) for $a\rightarrow 0$, which describes the non-spinning JMN-1 naked singularity spacetime.
Our purpose here to introduce Eq. (32) is that it could describe the spacetime of a slowly-spinning naked singularity, formed due to the transmutation of a spinning neutron star. For example, if the DM particles are accumulated inside a pulsar like the fastest known spinning pulsar PSR J1748-2446ad with $\omega\sim 716$ Hz [56] (i.e., $a_{\rm h}/M_{\rm h}\sim 0.09$), and it satisfies the collapse criterion, it could be transmuted into a slowly-spinning JMN-1 naked singularity of the same spin parameter (i.e., $a/M\sim 0.09$), provided that the dynamical ejecta is negligible. Even for a typical rapidly-spinning neutron star, the maximum spin parameter range could be around $a_{\rm h}/M_{\rm h}\sim 0.1$ (but see [57]). In such a scenario, one can study the evolution of the endoparasitic naked singularity and transmutation of a neutron star with the help of slowly-spinning JMN-1 naked singularity metric (Eq. 32) following [52] instead of Kerr spacetime. However, this metric is not enough to study the evolution of an endoparasitic naked singularity formed inside a fast-spinning white dwarf [53], earth-like planet, sun-like star etc., as the spin parameter is very high for these objects [53].

### v2

@@ 4.1 DM particles cloud modeled as a perfect fluid or inhomogeneous dust
$T^{\alpha}_{\beta}={\rm diag}\left(-\rho,p,p,p\right),$ (4.1)
@@
$\rho=\frac{F^{{}^{\prime}}}{R^{{}^{\prime}}R^{2}}$ (4.3)
@@
$R(t,r)=\left(r^{\frac{3}{2}}-\frac{3}{2}\sqrt{F(r)}t\right)^{\frac{2}{3}},$ (4.5)
@@
$t(r,s)=\frac{2r^{3/2}}{3\sqrt{F(r)}}\left(1-s^{3/2}\right)$ (4.6)
@@
$t_{v}(r)=\frac{2r^{\frac{3}{2}}}{3\sqrt{F(r)}}$ (4.7)
by substituting $R(t_{v},r)=0$ in eq. (4.5) or $s=0$ in eq. (4.6). This, in fact, gives the singularity curve which depicts the spacetime singularity formation as the collapse endstate. Note that the collapsing DM fluid spacetime (eq. 4.2) is matched smoothly with the exterior Schwarzschild spacetime. However, the singularity formed in the DM dust collapse scenario is event-like singularity. This means that the singularity initially becomes visible in principle, as the event horizon does not form. In a later stage, the singularity gets covered within the event horizon, provided that the energy conditions ensuring the positivity of mass and energy density are satisfied. Thus, the conclusion is, the formation of a tiny endoparasitic BH could not be avoided eventually for an inhomogeneous dust collapse. The similar situation could arise, even if the DM cloud is modeled as a perfect fluid, and, hence, the conclusion remains same.
@@ 4.2 DM particles cloud modeled as an anisotropic fluid
$ds^{2}=-e^{2\nu}dt^{2}+\frac{R^{\prime 2}}{H}dr^{2}+R^{2}d\Omega^{2}\;$ (4.8)
@@
Eqs. (4.8) and (4.9) reveal that one obtains six unknowns ($\rho,p_{\theta},\nu,H,R$ and the mass function $F$ for the DM cloud) that appear due to the integration of the Einstein equations. As there are four Einstein equations, we are free to choose two functions. The free functions are chosen to be the mass function $F(r)$, which is the initial mass that is conserved for the DM cloud, and the tangential pressure of it. Although $F(r)$ is shown to be independent of time [11], the tangential pressure depends on $r$ and $t$ via [11] $v(r,t)$ as $p_{\theta}\equiv p_{\theta}(r,v)$. A global choice for the pressure at all times fully fixes the evolution of the system.
@@
$F(r)=M_{\rm e}r^{3},\quad v_{\rm e}(r)=r^{\alpha},\quad F(R)=M_{\rm e}R^{\frac{3}{\alpha+1}},$ (4.10)
@@
$ds^{2}_{e}=-(1-M_{\rm e})\left(\frac{R}{R_{\rm b}}\right)^{\frac{M_{\rm e}}{1-M_{\rm e}}}dt^{2}+\frac{dR^{2}}{1-M_{\rm e}}+R^{2}d\Omega^{2}\;,$ (4.11)
@@
$\frac{F}{R_{\rm e}}=1-H_{\rm e}=\frac{4p_{\theta\rm e}}{\rho_{\rm e}+4p_{\theta\rm e}}\;.$ (4.12)
@@ 5 Transmutation timescale of the host neutron star
In section 4, we show that the endoparasitic collapsed object could be a BH (see section 4.1) or a naked singularity (see section 4.2) depending on the initial conditions. However, after its formation it will start to accrete matter immediately from its host neutron star with the spherical Bondi accretion rate:
$\frac{dm}{dt}\Bigg{|}_{\rm acc}=k_{1}m^{2}\,,$ (5.1)
@@
Figure 1: Mass ($M_{0}$) of the endoparasitic collapsed object (BH or naked singularity) formed inside a neutron star of $M_{\rm h}=1.4M_{\odot}$ at $t=t_{0}$ versus the mass ($m_{\chi}$) of DM particles. This figure explains that $M_{\rm h}>>M_{0}$. See section 5 for details.
Now, by integrating eq. (5.1):
@@
$\Rightarrow t_{\rm d}-t_{0}$ $=$ $\frac{1}{k_{1}M_{0}}\left[1-\frac{M_{0}}{M_{\rm h}}\right],$ (5.2)
@@
$\Rightarrow t_{\rm d}$$\approx$ $t_{0}+\frac{c_{\rm s}^{3}R_{\rm h}^{3}}{3G^{2}M_{\rm h}M_{0}}\equiv t_{0}+t_{\rm acc}$ (5.3)
where, $t_{\rm acc}$ is the accretion timescale after the formation of endoparasitic collapsed object. To obtain eq. (5.3) from eq. (5.2), we neglect the term $M_{0}/M_{\rm h}$ as $M_{\rm h}>>M_{0}$ which is clear from figure 1. The solid magenta and dashed purple curves of figure 1 indicate the mass ($M_{0}$) of the newly-formed endoparasitic collapsed object inside a typical neutron star of $M_{\rm h}=1.4M_{\odot}$ by collapsing of particles, respectively. Note that the Bondi (spherical) accretion rate (eq. 5.1) should be same [51] for a BH and a naked singularity. Since this should also be true for eq. (5.3), we only need to know $M_{\rm h}$ and the mass of the endoparasitic collapsed object $M_{0}$ (see eq. 3.7) to calculate $t_{\rm d}$, regardless of its nature.
@@
Figure 2: The variation of transmutation timescale ($t_{\rm d}$) of a neutron star of $M_{\rm h}=1.4M_{\odot}$ and $R_{\rm h}=10$ km as a function of the mass ($m_{\chi}$) of non-self-annihilating DM particles. Panels (a)-(b) correspond to two different DM density profiles with a fixed DM particle-nucleon scattering cross section, i.e., $\sigma>10^{-44}$ cm2. The black dotted horizontal line represents the Galactic age ($t_{\rm Ga}$). This figure shows that there is a non-zero possibility of transmutation of the neutron star into a collapsed object (BH or naked singularity depending on the initial conditions) in a reasonable time for a wide range of $m_{\chi}$ values, depending on the value of DM density profile ($\rho_{\rm DM}$). See section 5 for details.
@@
Figure 3: Constraints on the DM particle-neutron scattering cross section $\sigma$, as a function of mass ($m_{\chi}$) of the DM particles for the possible transmutation of a typical neutron star, if it is located at a distance $D$ with the specified $\rho_{\rm DM}$ value mentioned in the caption of each panel. The parameter space located on the right side of the red/blue-dashed curve indicates a transmutation for $t_{\rm d}<t_{\rm Ga}$. On the other hand, the left side of the red/blue-dashed curve suggests that the transmutation may occur in the future. See section 5 for details.
@@
For the purpose of demonstration to calculate $t_{\rm d}$, let us assume that a DM particle is captured after a single scattering, and, hence, we consider here the non-self-annihilating DM particles with a mass range: $1$ GeV $<m_{\chi}<10^{6}$ GeV [28]. If a typical neutron star of $M_{\rm h}\sim 1.4M_{\odot}$ and $R_{\rm h}\sim 10$ km captures the DM particles in the above mass range and forms an endoparasitic BH/naked singularity in its core, the entire neutron star could transmute into a BH/naked singularity depending on the initial condition of collapse as discussed in section 4. Now, if the transmutation time ($t_{\rm d}$) of a neutron star satisfies the following condition: $t_{\rm d}<t_{\rm Ga}$ (where $t_{\rm Ga}\sim 1.36\times 10^{10}$ yr is the Galactic age), the neutron star could have already transmuted. Note that $t_{\rm d}$ depends on the DM density ($\rho_{\rm DM}$) profile (see eq. 3.1, eq. 3.8 and eq. 5.3) as well. Thus, for the purpose of demonstration, here we consider the most commonly used Navarro-Frenk-White (NFW) DM density profile [52],
$\rho_{\rm DM}(D)$ $=$ $\rho_{0\rm DM}~{}(D/D_{s})^{-\alpha}(1+D/D_{s})^{-3+\alpha}$ (5.4)
@@
where, $D$ is the distance from the Galactic center. We consider the local DM density $\rho_{0\rm DM}=0.4$ GeV.cm-3 following [52, 47], the scale radius $D_{s}=20$ kpc following [33] with $\alpha=1$ [33]. Thus, $\rho_{\rm DM}\sim 7\times 10^{4}$ GeV.cm-3 [33] is obtained at a distance $D\sim 0.1$ pc inside the Galactic center, and $\rho_{\rm DM}\sim 0.4$ GeV.cm-3 is for the Galactic disk ($D\sim 10$ kpc) [53]. Substituting all the values in eq. (5.3), one can plot $t_{\rm d}$ versus $m_{\chi}$ for the DM particle-neutron cross section $\sigma>10^{-44}$ cm-2 for the two different values of DM density.
@@
Panel (a) of figure 2 shows that the transmutation of a neutron star could not be possible in the Galactic disk for the fermionic DM particles, whereas it could be possible for the bosonic DM particles. Panel (b) corresponds to the Galactic bulge, and it shows that the transmutation could be possible for both the fermionic and bosonic DM particles. However, comparing both the panels we can see that the possibility of transmutation within $t_{\rm d}\leq t_{\rm Ga}$ decreases with decrement of $\rho_{\rm DM}$. Note that figure 2 stands for $\sigma>10^{-44}$ cm-2 only, and, thus it can be considered just as an example of how $t_{\rm d}$ varies with $m_{\chi}$ for three different $\rho_{\rm DM}$ values with a fixed value of $\sigma$. One cannot constrain the $\sigma-m_{\chi}$ parameter space from figure 2. Panels (a)-(b) of figure 3, in fact, help to constrain the non-self-annihilating DM-neutron scattering cross section $\sigma$, as a function of $m_{\chi}$ for the possible transmutation within $t_{\rm d}\leq t_{\rm Ga}$, if the neutron star is located in the Galactic disk and Galactic bulge, respectively. The parameter space located on the right side of the red (blue-dashed) curve indicates a transmutation for $t_{\rm d}\leq t_{\rm Ga}$. On the other hand, the left side of the red (blue-dashed) curve suggests that the transmutation may occur in the future.
@@ 6 End state of the transmuted object
Now, the question is, what would be the end state of the transmuted object. The tiny endoparasitic BH formed (as discussed in section 4.1) due to the collapse of a perfect DM fluid or dust should transmute the host into a BH. The transmutation process should be as discussed in various papers (e.g., [27, 29]). If the host neutron star has a non-zero spin (e.g., a pulsar), the transmutation process could be evolved as described in [54]. However, the end state of the transmuted object would be a BH, and not a naked singularity as shown in [55].
When an endoparasitic naked singularity (as discussed in section 4.2) is formed due to an anisotropic DM fluid collapse, it starts to accrete matter from its host and eventually the entire host would transmute into a naked singularity. As an exactly non-spinning object may not exist in the Universe [56], the host neutron star should have a non-zero spin parameter $a_{\rm h}$ $\equiv J_{\rm h}/M_{\rm h}$, where $J_{\rm h}$ is the angular momentum of the host. Since the spin frequency ($\omega$) of both the neutron star and the endoparasitic naked singularity during its formation (i.e., at ${t\rightarrow t_{0}}$) should be the same (as $M_{0}<<M_{\rm h}$) [55], the endoparasitic naked singularity should also have a non-zero spin parameter. Let us denote that spin parameter of the endoparasitic naked singularity as $a_{0}$ at time $t\rightarrow t_{0}$. However, if one calculates the numerical value of $a_{0}/M_{0}$, it comes as negligible small (as $M_{0}<<M_{\rm h}$), and, hence, one can consider it as $a_{0}\rightarrow 0$ for all the practical purposes. During accretion from its host, the value of spin paramater starts to increase from $a_{0}$ and could eventually acquire $\sim a_{\rm h}$ [55] at the end of the transmutation process (i.e., $t=t_{\rm d}$), if the dynamical ejecta is small [54, 55]. If the dynamical ejecta is large [55] during accretion by the endoparasitic naked singularity, the final transmutation could give birth to a sub-solar-mass or near-solar mass slowly-spinning JMN-1 naked singularity. Computation of the evolution of such an endoparasitic naked singularity, which is a complex and entirely unexplored field, will be considered in future research works.
@@
$-2a\sin^{2}\theta\left[\left(\frac{R}{R_{\rm b}}\right)^{\frac{M_{\rm e}}{2(1-M_{\rm e})}}-(1-M_{\rm e})\left(\frac{R}{R_{\rm b}}\right)^{\frac{M_{\rm e}}{(1-M_{\rm e})}}\right]dtd\phi$ (6.1)
where, $a$ is the spin parameter of this spacetime. The above slowly-spinning JMN-1 spacetime (eq. 6.1) can be matched with the exterior slowly-spinning Kerr spacetime at $R=R_{\rm b}$ [57],
@@
where $2M=M_{\rm e}R_{\rm b}$ is the Schwarzschild mass of the slowly-spinning ($a/M<<1$) naked singularity. Eq. (6.1) reduces to eq. (4.11) for $a\rightarrow 0$, which describes the non-spinning JMN-1 naked singularity spacetime.
Our purpose here to introduce eq. (6.1) is that it could describe the spacetime of a slowly-spinning naked singularity, formed due to the transmutation of a spinning neutron star. For example, if the DM particles are accumulated inside a pulsar like the fastest known spinning pulsar PSR J1748-2446ad with $\omega\sim 716$ Hz [58] (i.e., $a_{\rm h}/M_{\rm h}\sim 0.09$), and it satisfies the collapse criterion, it could be transmuted into a slowly-spinning JMN-1 naked singularity of the same spin parameter (i.e., $a/M\sim 0.09$), provided that the dynamical ejecta is negligible. Even for a typical rapidly-spinning neutron star, the maximum spin parameter range could be around $a_{\rm h}/M_{\rm h}\sim 0.1$ (but see [59]). In such a scenario, one can study the evolution of the endoparasitic naked singularity and transmutation of a neutron star with the help of slowly-spinning JMN-1 naked singularity metric (eq. 6.1) following [54] instead of Kerr spacetime. However, this metric is not enough to study the evolution of an endoparasitic naked singularity formed inside a fast-spinning white dwarf [55], earth-like planet, sun-like star etc., as the spin parameter is very high for these objects [55].